A sample of gas in a container has a volume of 18.10mL at 3.500 atm. What will be the volume if the pressure becomes 2.500 atm, with a fixed amount of gas and temperature? 


1
Expert's answer
2021-05-13T04:57:13-0400

Initial Volume, "V_1=18.10\space mL"

Initial pressure, "P_1=3.500\space atm"

Final Volume "=V_2"

Final pressure, "P_2=2.500\space atm"

At constant temperature and mass

"P_1V_1=P_2V_2"

"V_2=V_1\dfrac{P_1}{P_2}=18.10\times\dfrac{3.5}{2.5}=25.34\space mL"

"V_2=25.34\space mL"